In how many ways can 6 students be seated in a row of 6 seats if 2 of the students insist on sitting beside each other?
Answer by ikleyn(52782)

.
In how many ways can 6 students be seated in a row of 6 seats
if 2 of the students insist on sitting beside each other?
~~~~~~~~~~~~~~~~~~~~~

First, we consider this inseparable couple as one object.

Doing this way, we have 5 objects to order in a row, and
we can order 5 objects in 5! = 5*4*3*2*1 = 120 different ways.


But we also can order this pair in two different ways - it gives us
the total of 2*120 = 240 different possible ordering
of 6 students under given condition.

Solved, with complete explanations.